OpenStudy (anonymous):

Help please someone~~ The outline of a circularly symmetric bottle has a height of 1 metre. The outline fits the curve x^2=(ky(1-y)^4+0.1)^2 where x and y are distances in metres, k>0 and 0 less than or equal to x less than or equal to 1 state the integral used to determine the volume of the bottle Find k when the volume = 100 litres ( 1 litre = 0.001 m^3 )

OpenStudy (anonymous):

o.. dont i need a pi?

OpenStudy (anonymous):

i thought it has something to do with volume of revolution

OpenStudy (anonymous):

then it asks for find k i integrate from what to what?

OpenStudy (anonymous):

V=pi\[\int\limits_{a}^{b?}x^2dy\]

OpenStudy (anonymous):

v=volume so a=0 and b=1 get that

OpenStudy (anonymous):

yeap i know that

OpenStudy (anonymous):

so when volume =100 i solve for k... but i got a weird answer

OpenStudy (anonymous):

the second question asks me to evaluate this integral interms of k

OpenStudy (anonymous):

when i evaluated it using calculator its like a whole bunch of long numbers and its a quadratic function

OpenStudy (anonymous):

now try to multiply the right hand side of ur equation u'll get ...x^2=k(y^5-4y^4+6y^3-4y2+y) this is what u need to integrate in order to find k are good with this?

OpenStudy (anonymous):

x^2=(ky(1-y)^4+0.1)^2 why do i need multiply my right hand side equation? multiply with what...? there is alrdy a k in there mind if u could show full working out? thx.. if u can

OpenStudy (anonymous):

okay! i made a mistake with my multiplication. let us work out what is in the first bracket 0.1=0 so will get x^2=(ky(1-y)^4)^2 get that

OpenStudy (anonymous):

how is 0.1 equals 0 ..... ?

OpenStudy (anonymous):

is that zero or the letter o?

OpenStudy (anonymous):

its zero lol

OpenStudy (anonymous):

hahaha so zero times any number is isn't it zero? lol

OpenStudy (anonymous):

yea... but why is 0.1 = 0..? it doesnt make sense in thefirst place

OpenStudy (anonymous):

oh! is it 0,1(comma) or 0.1(times)?

OpenStudy (anonymous):

what...? x^2=(ky(1-y)^4+0.1)^2 its + 0.1

OpenStudy (anonymous):

i'm confused about thsat last expression is it 0comma 1 or 0 times 1?

OpenStudy (anonymous):

0.1 its 0 point 1

OpenStudy (anonymous):

if u have a smart calculator multiply it out and give me the expression without the brackets so that we can start from that

OpenStudy (anonymous):

dnt have one lol..

OpenStudy (anonymous):

any way God granted us a best tool our heads let use them and find what that is okay?

OpenStudy (anonymous):

haha ok..

OpenStudy (anonymous):

did u get it?

OpenStudy (anonymous):

nope..

OpenStudy (anonymous):

let do it this way x2=[ky(1-y)4 +0.1] [ky(1-y)4 + o,1] get this ?

OpenStudy (anonymous):

wow after multiplying out i found :.... x^2=k^2[y^10-8y^9+28y^8-52y^7+72y^6-64y^5+40y^4-16y^3+3y^2] +0.01

OpenStudy (anonymous):

K=0,11183 is this the correct answer